OpenStudy (anonymous):

what is the slope of the equation 11x-y=-3?=11

OpenStudy (anonymous):

is the ques. 11x-y=-3=11 in the right way?!

OpenStudy (anonymous):

yah!!

OpenStudy (anonymous):

well the slope is 11

Directrix (directrix):

11x-y=-3?=11 Please check the equation. It should not have two = symbols in it. Also, 3? is not a number.

OpenStudy (anonymous):

11x-y=-3

OpenStudy (anonymous):

in any equation y=mx+b where m is the Slope n don't look on any other symbols

Directrix (directrix):

slope is 11

OpenStudy (anonymous):

Coefficient of x= slope Slope=11